A rental car costs a minimum charge plus a fixed charge per mile driven. one customer was charged $60 for 160 miles of use. Another customer drove 280 miles and paid $90. Write an equation relating the total cost, C(in dollars), to the miles driven,x. Please show all work thanks
Answer by stanbon(75887)

A rental car costs a minimum charge plus a fixed charge per mile driven. one customer was charged $60 for 160 miles of use. Another customer drove 280 miles and paid $90. Write an equation relating the total cost, C(in dollars), to the miles driven,x. Please show all work thanks
-----------------------------------------------
You have two points: (160,60) and (280,90)
------------
slope: (90-60)/(280-160) = 30/120 = 1/4
-------------------
intercept:
60 = (1/4)(160) + b
b = 60 - 40 = 20
=====================
EQUATION:
C(m) = (1/4)m + 20 where m is the distance in miles and C is the cost.
=======================================================
Cheers,
Stan H.